A particle moving with uniform acceleration in a straight line was first observed to be moving at 4.5m/s. After 10 seconds, it was moving at 6.8m/s. Find its acceleration.
1
Expert's answer
2020-10-05T10:56:26-0400

The acceleration is


"a=\frac{\Delta v}{\Delta t}=\frac{6.8-4.5}{10}=0.23\text{ m/s}^2."